OpenStudy (anonymous):

The function f(x) = 2(3)x represents the growth of a dragonfly population every year in a remote swamp. Rose wants to manipulate the formula to an equivalent form that calculates seven times a year, not just once a year. Which function is correct for Rose's purpose, and what is the new growth rate?

OpenStudy (help!!!!):

so 7 years? f(x)=2(3)*7x

OpenStudy (anonymous):

its saying 7 times in one year i think

OpenStudy (help!!!!):

oh then just plug in 7 for x

OpenStudy (anonymous):

really ?wow

OpenStudy (anonymous):

f(x) = 2(3)x; growth rate 200% f(x) = 2(1.17)7x; growth rate 17% f(x) = 2(3)x, growth rate 2% f(x) = 2(1.17)x; growth rate 2% thats what it gives me thow

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (anonymous):

@jdoe0001